Given P(A)=0.22, P(B)=0.72, and P(A∩B)=0.18, what is P((A∩B)^c)?
Answer by ikleyn(52944)

.

P((A ∩ B)^c) is the complement to  P((A ∩ B)), and, as such, is equal to


    1 - P((A ∩ B)) = 1 - 0.18 = 0.82.


The rest of the input data in the post is IRRELEVANT.